OpenStudy (anonymous):

Represent the complex number graphically, and find the trigonometric form of the number. -2 - 2i

OpenStudy (anonymous):

This is what I have so far: The result will be in Quadrant III. |a + bi| = sqrt[a2+ b2] = sqrt[(-2)2+ (-2)2] = sqrt[4+ 4] = sqrt[8] = 2sqrt[2] r = 2sqrt[2] tanthetar = |b/a| theta= [(-2)/(-2)] tantheta= |1| theta = tan-1 (1) theta = 45 degrees Trigonometric form: 2sqrt[2] [(cos45degrees) + (isin45degrees)]

OpenStudy (anonymous):

My teacher said to find a reference angle to 45 degrees. Would I add 180 to 45 to achieve the answer?